OpenStudy (anonymous):

Find the total resistance of a 4Ώ and a 5Ώ resistor that are wired in parallel. please show work

OpenStudy (anonymous):

try the physics openstudy maybe?

OpenStudy (maya):

(1/4 + 1/5)^-1

OpenStudy (maya):

This will give you 20/9 = total resistance

OpenStudy (saifoo.khan):

.45

OpenStudy (saifoo.khan):

9/20 = 0.45

OpenStudy (maya):

2.22

OpenStudy (saifoo.khan):

nooo, it will b 0.45

OpenStudy (anonymous):

i think maya is right you want the inverse of the sum of the inverses

OpenStudy (anonymous):

2.22

OpenStudy (maya):

right

OpenStudy (anonymous):

maya is correct 2.22222222222 and so on

OpenStudy (maya):

:)

OpenStudy (saifoo.khan):

ohh sorrry!

OpenStudy (saifoo.khan):

Yeah, she is crrect.. really sorry! :/

OpenStudy (saifoo.khan):

crrct maya. (:

OpenStudy (anonymous):

In parallel resistors combine as: \[\frac{1}{R_1}+\frac{1}{R_2}+...+\frac{1}{R_n}=\frac{1}{R_{total}}\]

OpenStudy (anonymous):

I only have 7 medals and i really need 14 :(

OpenStudy (saifoo.khan):

i just hav 24.. I need 48. :/ Jk. ;)

OpenStudy (anonymous):

helpp if you tell me your favorite equation ill give you a medal

OpenStudy (anonymous):

of anything?

OpenStudy (anonymous):

yep any equation

OpenStudy (anonymous):

I got one vitale!!!! F=ma :DDD I'm a physics major >.>

OpenStudy (anonymous):

ok haha

OpenStudy (anonymous):

C=(pi symbol)d

OpenStudy (anonymous):

k that works

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

yeah i clicked it

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

8 more to go!

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

anyone else wanna give me one?

OpenStudy (saifoo.khan):

wht u gonna do with thm??

OpenStudy (anonymous):

nothinn just want them :)

OpenStudy (saifoo.khan):

why not try the "exchange thing"??

OpenStudy (anonymous):

k medal for medal?

OpenStudy (saifoo.khan):

hehe, Yeah.

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

Thanks <3

OpenStudy (saifoo.khan):

ui got it. ;) <3

OpenStudy (anonymous):

anyone else wanna join?

OpenStudy (saifoo.khan):

U*